# A statistical test for network similarity

### v1

@@ Abstract
In this article, we revisit and expand our prior work on graph similarity. In this version of our work, we offer an extensive array of empirical tests. We also examine the sensitivity of our test to network variations. Our test performs exactly as expected, on synthetic and real-world graphs. It offers a very accurate measure of graph (dis)similarity.
@@ 1 Introduction
In this article, we revisit and expand our prior work on graph similarity [18]. As with our earlier work, we focus on a view of similarity which does not require node correspondence between graphs under comparison. Our work is suited to the temporal study of networks, change-point and anomaly detection and simple comparisons of static graphs. This version of our work includes extensive empirical testing and a detailed sensitivity analysis. In addition, it also includes an extension of our test to directed graphs.
The goal of this work is to offer a metric to compare and assess the (dis)similarity of two networks. Given two graphs, $G_{1}$ and $G_{2}$, we seek to answer the question “how different is $G_{1}$ from $G_{2}$?”. Furthermore, because we propose a metric, given a graph $G_{3}$, our metric also offers an answer to the question “is $G_{3}$ more like $G_{1}$ or $G_{2}$?”.
Graph isomorphism is a problem for which there is no known polynomial-time solution. The more general problem of computing graph similarity metrics, graph edit distance or maximum common subgraph, is NP-hard. Nevertheless, with the widespread availability of graph data sets (social networks, communication networks, financial networks,… ) comparing networks became and remains a very important task [30, 35, 23].
To meet this need to compare graphs, several methods have been created [29]. Our method stands out from others, because of its statistical grounding, its integration of node-level, neighborhood-level and graph-wide features and polynomial time complexity.
@@
Above all, Jaccard distance can be interpreted through a probabilistic lens. It is because of this interpretation that we are then able to compare networks as probability distributions of node-node distances, by applying common statistical techniques. This conversion to a distance matrix and its probabilistic examination as a distribution offers a broad structural viewpoint of the graph.
@@ 2 Previous work
Before discussing the work which inspired and guided us, we begin this section on prior work with an area of the literature which we deliberately avoided, graphons. While we are aware of this active area of the literature, we do not consider it well-suited to the objectives of our research.
2.1 Graphons
@@ 2.2 Work we considered
Our work is heavily influenced by that of Schieber et al. (2017) [27]. Like these authors, we study the distribution of node-node distances. We are also guided in our work by Wang et al. (2022) [33]. Like these authors, we also consider node to node distance. Nevertheless, we do not compute distances between vector representations of nodes. Unlike these authors and in light of the limitations (of shortest path distances) identified by Akara et al.  (2021) [2] and the work of Camby and Caporossi (2017) [5], we use Jaccard distance between nodes. The details and benefits of using Jaccard for node similarity have been well-documented [12, 5, 21, 20] and so have the shortcomings of shortest-path [6, 9].
@@ 3 Methods
Our (dis)similarity algorithm involves three distinct steps. First, we convert each graph into an all-pairs (Jaccard) distance. We then interpret these distances matrices as sets of random variable. Finally, we measure the (dis)similarity between each graph (random variable set) using the Kolmogorov-Smirnov metric [1, 25].
@@ 3.1 All-pairs distance
Our analysis begins with a transformation of each graph being compared into an all-pairs (Jaccard) distance matrix. For each pair of vertices in a given graph we compute the following distances.
@@
As mentioned earlier, in this version of our work, we extend this distance to directed graphs. In the case of directed graphs, we simply decompose the cardinalities of the intersection and union of shared connections. We decompose those cardinalities into the sum of the cardinality of their two constituents:
$\zeta_{ij}=1-\underbrace{\frac{|p_{i}\cap p_{j}|+|s_{i}\cap s_{j}|}{|p_{i}\cup p_{j}|+|s_{i}\cup s_{j}|}}_{\sigma_{ij}}\in[0,1]\,.$
In this specific case, $p_{i}\,(p_{j})$ is the set of predecessor nodes to node i(j)i\,(j). Meanwhile, $s_{i}\,(s_{j})$ is the set of successor nodes.
@@ 3.1.1 Probabilistic interpretation of the Jaccard distance
In the general (and undirected) case, Jaccard similarity ($\sigma_{ij}$) between two nodes $i$ and $j$ has a probabilistic interpretation. Consider all nodes of a network excluding $i$ and $j$ and select at random a node $k$. Here, Jaccard similarity is an estimate of the (conditional) probability that both $i$ and $j$ are connected to $k$, given that at least one of $i$ or $j$ is connected to $k$. Mathematically, we express $\sigma_{ij}$ as
$\sigma_{ij}=P\left(\left(e_{ik}\land e_{jk}\right)\mid\left(e_{ik}\lor e_{jk)}\right)\right)\,$,
@@
Naturally, Jaccard distance is its complement (i.e., $\zeta_{ij}=1-\sigma_{ij}$). Therefore, it can be interpreted as one of these two cases,
@@
    the (conditional) probability that $i$ is connected to $k$, but $j$ is not,
@@
    the (conditional) probability that $j$ is connected to $k$, but $i$ is not.
@@
$\zeta_{ij}=1-\sigma_{ij}=1-P\left(\left(e_{ik}\land e_{jk}\right)\mid\left(e_{ik}\lor e_{jk)}\right)\right)=P\left(\left(e_{ik}\veebar e_{jk}\right)\mid\left(e_{ik}\lor e_{jk)}\right)\right)\,.$
In the specific case of directed graphs, the $\zeta_{ij}$ has an essential similar interpretation.
@@ 3.1.2 From graph to empirical probability distribution
With all distances $\zeta_{ij}$ computed, we then examine their statistical distribution. In light of the probabilistic interpretation of the $\zeta_{ij}$, we treat these quantities as random variables. We can therefore compare graphs as empirical probability distributions of node-node distances.
Figure 1 illustrates the core ideas of this work. On the left, we see the distribution of node-node distances for an Erdős-Rényi (ER) graph [8, 10]with 5,000 nodes and an edge probability of $p=0.5$. On the right, we see the distribution of distances between nodes of a planted partition model graph (PPM) [7] with 50 blocks of 100 nodes and in/out edge probabilities of $0.9/0.1$. (Both graphs have the same number of nodes, both distributions contain the same number of observations.)
@@
Figure 1: Distances as distributions
We immediately note that the very significant structural difference between these two graphs is very well translated into the difference between the distributions of node-node distances. The ER graph’s distances are symmetrically distributed about their mean, in a Gaussian-like pattern. Markedly, the PPM graph’s distances are left-skewed and bi-modal. The left mode reflects distances between nodes in the same blocks, whereas the right mode reflects distances between nodes not in the same blocks. Naturally, this bi-modal pattern does not appear in the case of the ER graph.
@@
Unfortunately, the K-S test requires that both the observations within each distribution and the distributions being compared themselves be independent and identically distributed (iid). While the p-values in our test do behave exactly as expected, they cannot be relied on for their statistical accuracy. For this reason, we focus our attention on the K-S distances themselves. After all, even with moderately sized graphs like the one in our experiments (e.g., $N=5,000\Rightarrow\sim 5,000^{2}$ distances), the sizes of our samples are extremely large. Argubaly, these extremely large samples make our distances very reliable metrics of the (dis)similarity between the distributions being compared.
@@ 4 Numerical results
We conduct three types of numerical experiments. In Section 4.1, we compare graphs that are known to be different. The goal of these comparisons is to validate the correct rejection of the null hypothesis. These tests are new instances of the tests we conducted in our previous work [18].
In Section 4.2, we compare individual graphs with versions of themselves with randomly selected nodes or edges having been removed. The goal of these experiments is, in part, to validate the correct non-rejection of the null, in cases where we compare the graph to an unaltered version of itself. These trials also offer insight into the sensitivity of our test to graph structure modification through node or edge removal (e.g., in cases of temporal or change-point detection settings).
@@ 4.1 Two-graph tests
As stated earlier, the goal of these comparisons is to validate the correct rejection of the null hypothesis. We conduct 12 tests between synthetic graphs with known (dissimilar) structures. Table 1 shows cases where the structural differences between the graphs are arguably slight. The distance distributions for each of the graphs referenced in this table are presented in Figure 2 (Erdös-Rényi, ER) and Figure 3 (stochastic block model, SBM).
@@
All synthetic graphs in these experiments have the same number of vertices ($N=3,400$). The SBM graphs are all composed of 45 blocks of varying sizes ($n\in[50,99]$) but uniform in/out edge probabilities. These probabilities are shown in their naming convention. The ER graphs against which they are compared were generated using edge probabilities that are equal to the density of the SBM graphs in the comparisons. The CM graphs were generated using a power law in/out degree sequences of exponent of 3.5.
Table 1: Pairwise comparisons of ER and SBM graphs. (hard cases)
Graph 1 Graph 2 KS Statistic p-value
ER 0.309 SBM 0.7/0.3 0.004 0
ER 0.309 (Directed) SBM 0.7/0.3 (Directed) 0.005 0
ER 0.213 SBM 0.8/0.2 0.013 0
ER 0.213 (Directed) SBM 0.8/0.2 (Directed) 0.018 0
ER 0.118 SBM 0.9/0.1 0.061 0
ER 0.118 (Directed) SBM 0.9/0.1 (Directed) 0.086 0
Table 2: Pairwise comparisons of CM, ER, and SBM graphs. (easier cases)
Graph 1 Graph 2 KS Statistic p-value
CM3400 ER 0.213 0.9988 0
CM3400 (Directed) ER 0.213 (Directed) 0.9982 0
CM3400 SBM 0.9/0.1 0.9988 0
CM3400 (Directed) SBM 0.9/0.1 (Directed) 0.9977 0
SBM 0.9/0.1 ER 0.213 0.9762 0
SBM 0.9/0.1 (Directed) ER 0.213 (Directed) 0.9774 0
@@
Figure 2: Jaccard distances distributions for ER models.
@@
Figure 3: Jaccard distances distributions for stochastic block models.
@@
Figure 4: Jaccard distances distributions for configuration models.
@@
4.2 Sensitivity analysis and real-world graphs
The tests in this section have three different goals. First we compare each graph to itself. Here, we expect the null not be rejected. We also compare each graph to a version of itself with a portion of its nodes or edges removed. The goal of these trials is to examine the test’s sensitivity to various structural modifications. We compare every graph in this section to a version of itself with 0, 0.5, 1, 5, 10, 25% of its (randomly selected) nodes or edges removed.
@@ 4.2.1 Synthetic graphs
We begin by removing a percentage of randomly selected nodes from each of our four ER graphs, and compare the distribution of Jaccard distances to the distribution of distances in the complete original graph. The distance D is shown, while the corresponding test p-value is presented in parentheses.
@@
Percent Removed 0 0.5 1 5 10 25
ER 0.5 0 (1) 0.0011 (0.0012) 0.0029 (9e-22) 0.0084 (8e-170) 0.0134 (0) 0.0382 (0)
ER 0.5 (Directed) 0 (1) 0.0013 (6e-05) 0.0023 (5e-14) 0.0073 (1e-127) 0.0180 (0) 0.0387 (0)
ER 0.333 0 (1) 0.00085 (0.0307) 0.0030 (5e-22) 0.0080 (3e-152) 0.0158 (0) 0.0383 (0)
ER 0.333 (Directed) 0 (1) 0.00077 (0.0652) 0.0019 (9e-10) 0.0070 (2e-117) 0.0135 (0) 0.0408 (0)
@@
Percent Removed 0 0.5 1 5 10 25
ER 0.5 0 (1) 0.0953 (0) 0.1886 (0) 0.7586 (0) 0.9788 (0) 1 (0)
ER 0.5 Directed 0 (1) 0.1337 (0) 0.2629 (0) 0.9022 (0) 0.9989 (0) 1 (0)
ER 0.333 0 (1) 0.0522 (0) 0.1037 (0) 0.4810 (0) 0.7987 (0) 0.998 (0)
ER 0.333 Directed 0 (1) 0.0732 (0) 0.1457 (0) 0.6376 (0) 0.9290 (0) 1 (0)
@@
Percent Removed 0 0.5 1 5 10 25
SBM 0.7/0.3 0 (1) 0.002 (1e-9) 0.003 (1e-17) 0.006 (1e-95) 0.014 (0) 0.041 (0)
SBM 0.7/0.3 (Directed) 0 (1) 0.001 (0.129) 0.002 (1e-10) 0.007 (1e-114) 0.013 (0) 0.042 (0)
SBM 0.9/0.1 0 (1) 0.0009 (0.029) 0.002 (1e-10) 0.007 (1e-116) 0.016 (0) 0.036 (0)
SBM 0.9/0.1 (Directed) 0 (1) 0.0009 (0.019) 0.001 (1e-6) 0.008 (1e-180) 0.014 (0) 0.035 (0)
@@
Percent Removed 0 0.5 1 5 10 25
SBM 0.7/0.3 0 (1) 0.047 (0) 0.093 (0) 0.438 (0) 0.751 (0) 0.995 (0)
SBM 0.7/0.3 (Directed) 0 (1) 0.065 (0) 0.13 (0) 0.584 (0) 0.893 (0) 1 (0)
SBM 0.9/0.1 0 (1) 0.015 (0) 0.029 (0) 0.146 (0) 0.284 (0) 0.628 (0)
SBM 0.9/0.1 (Directed) 0 (1) 0.02 (0) 0.041 (0) 0.202 (0) 0.389 (0) 0.785 (0)
Throughout these experiments, our test behaves exactly as expected. The test detects dissimilarity as it increases and never classifies a graph as different from its unmodified self. We also note that, predictably, edge removal introduces a more marked variation in the graph than node removal.

### v2

@@ Abstract
In this article, we revisit and expand our prior work on graph similarity. As with our earlier work, we focus on a view of similarity which does not require node correspondence between graphs under comparison. Our work is suited to the temporal study of networks, change-point and anomaly detection and simple comparisons of static graphs. It provides a similarity metric for the study of (weakly) connected graphs. Our work proposes a metric designed to compare networks and assess the (dis)similarity between them. For example, given three different graphs with possibly different numbers of nodes, $G_{1}$, $G_{2}$ and $G_{3}$, we aim to answer two questions: a) ”How different is $G_{1}$ from $G_{2}$?” and b) ”Is graph $G_{3}$ more similar to $G_{1}$ or to $G_{2}$?”. We illustrate the value of our test and its accuracy through several new experiments, using synthetic and real-world graphs.
@@ 1 Introduction
In this article, we revisit and expand our prior work on graph similarity [18]. As with our earlier work, we focus on a view of similarity which does not require node correspondence between graphs under comparison. Our work is suited to the temporal study of networks, change-point and anomaly detection and simple comparisons of static graphs. Naturally, change-point or anomaly detection can be performed with our test, but it should be coupled with area expertise and a thorough examination of each graph’s structure. Our work provides a similarity metric. However, the significance of this (dis)similarity is more a structural and domain question than a statistical one.
While our earlier work did not explicitly mention this applicability condition of (weak) connectedness, we wish to clearly state it here. In this updated version, we show extensive empirical testing and a detailed sensitivity analysis. In addition, we also extend our test to (weakly connected) directed graphs. Our work proposes a metric designed to compare networks and assess the (dis)similarity between them. For example, given three different graphs with possibly different numbers of nodes, $G_{1}$, $G_{2}$ and $G_{3}$, we aim to answer two questions: a) ”How different is $G_{1}$ from $G_{2}$?” and b) ”Is graph $G_{3}$ more similar to $G_{1}$ or to $G_{2}$?”.
To this day, graph isomorphism remains a challenging problem, with no known polynomial-time solution. Meanwhile, the broader tasks of calculating graph similarity metrics, such as graph edit distance or maximum common subgraph, are NP-hard. However, with the growing availability of graph datasets, from social networks and communication networks to financial systems, network comparison and analysis have become crucial [31, 36, 24].
To meet this need to compare graphs, several methods have been created [30]. Our method stands out from others because of its statistical grounding, its integration of node-level, neighborhood-level and graph-wide features and polynomial time complexity.
@@
Above all, Jaccard distance can be interpreted through a probabilistic lens. It is because of this interpretation that we are then able to compare networks as probability distributions of node-node distances, by applying common statistical techniques. This conversion to a distance matrix and its probabilistic examination as a distribution offers a broader structural viewpoint of the graph than the focused examination of specific graph attributes.
@@ 2 Previous work
In our discussion of the work done to this point, we would like to begin with a justification for what we deliberately omitted, a discussion of graphons. While we are aware of this very active area of the literature, we do not consider it well-suited to the objectives of our research. Nevertheless, we feel it is necessary to reassure readers that graphons are not competitors for the solution we propose in this work.
@@
Our work is heavily influenced by that of Schieber et al. (2017) [28]. Following their example, we study the distribution of node-node distances. However, while these authors study the distribution of shortest-path distance between nodes, we use Jaccard. We are also guided by the work of Wang et al. (2022) [34]. Like Wang et al., we also examine the distribution of node to node distance, but with more clearly obtained distances. In contrast to them, we do not compute distances between vector representations of nodes that are obtained by random walk distances [23]. Unlike the work described in these two publications, in light of the overhead imposed by random walk simulations, the limitations of random walk [32, 33, 25, 18] or shortest path distances [6, 9, 2] and the work of Camby and Caporossi (2017) [5], we use Jaccard distance between nodes. In fact, the details and benefits of using Jaccard for node similarity have been well-documented [12, 5, 21, 20].
@@ 3 Methods
Our (dis)similarity algorithm involves three distinct steps. First, we convert each (weakly) connected graph into an all-pairs Jaccard distance. We then interpret these distances matrices as sets of random variables. Finally, we measure the (dis)similarity between each graph (random variable set) using the Kolmogorov-Smirnov metric [1, 26].
@@ 3.1 All-pairs distance
As described in our earlier work [18], our analysis begins with a transformation of each graph being compared into an all-pairs Jaccard distance matrix. For each pair of vertices in a given graph we compute the following distances.
@@
As mentioned earlier, in this version of our work, we extend this distance to directed graphs. In the case of directed graphs, we simply decompose the cardinalities of the intersection and union of shared connections into the sum of the cardinality of their two constituents:
$\zeta_{ij}=1-\underbrace{\frac{|p_{i}\cap p_{j}|+|\sigma_{i}\cap\sigma_{j}|}{|p_{i}\cup p_{j}|+|\sigma_{i}\cup\sigma_{j}|}}_{s_{ij}}\in[0,1]\,.$
In this specific case, $p_{i}\,(p_{j})$ is the set of predecessor nodes to node i(j)i\,(j). Meanwhile, $\sigma_{i}\,(\sigma_{j})$ is the set of successor nodes.
@@ 3.1.1 Probabilistic interpretation of the Jaccard distance
As outlined in our previous work, Jaccard similarity ($s_{ij}$) between two nodes $i$ and $j$ has a probabilistic interpretation, in the undirected case [18]. Indeed, Jaccard similarity is an estimate of the conditional probability that both $i$ and $j$ are connected to $k$, given that at least one of $i$ or $j$ is connected to $k$. Mathematically, we expressed $s_{ij}$ as
$s_{ij}=P\left(\left(e_{ik}\land e_{jk}\right)\mid\left(e_{ik}\lor e_{jk)}\right)\right)\,$,
@@
Naturally, Jaccard distance is its complement (i.e., $\zeta_{ij}=1-s_{ij}$). Therefore, it can be interpreted as one of these two cases,
@@
    the conditional probability that $i$ is connected to $k$, but $j$ is not,
@@
    the conditional probability that $j$ is connected to $k$, but $i$ is not.
@@
$\zeta_{ij}=1-s_{ij}=1-P\left(\left(e_{ik}\land e_{jk}\right)\mid\left(e_{ik}\lor e_{jk)}\right)\right)=P\left(\left(e_{ik}\veebar e_{jk}\right)\mid\left(e_{ik}\lor e_{jk)}\right)\right)\,.$
In the specific case of directed graphs, the $\zeta_{ij}$ has an essentially similar interpretation, but with edge direction ignored.
@@ 3.1.2 From graph to empirical probability distribution
With all distances $\zeta_{ij}$ computed, we then take a graph-wide point of view. In light of the probabilistic interpretation of the $\zeta_{ij}$ described previously, we compare graphs as empirical probability distributions of node-node distances. In other words, each node-node distance is understood to be a random variable and every graph is examined as an empirical statistical distribution.
Figure 1, borrowed from our previous work on the subject [18], provides an example of the application of the core ideas of this work. On the left, we see the distribution of node-node distances for an Erdős-Rényi (ER) graph [8, 10] with 5,000 nodes and an edge probability of $p=0.5$. On the right, we see the distribution of distances between nodes of a planted partition model graph (PPM) [7] with 50 blocks of 100 nodes and in/out edge probabilities of $0.9/0.1$. (Both graphs have the same number of nodes, both distributions contain the same number of observations.)
@@
Figure 1: Distances as distributions. The distribution of node-to-node distances for the Erdős-Rényi (ER) graph is unimodal while it is bimodal for the planted partition model (PPM) graph.
Here, it is clear that the very significant structural difference between these two graphs is very well translated into the difference between the distributions of node-node distances. The ER graph’s distances are symmetrically distributed about their mean, in a Gaussian-like pattern. Markedly, the PPM graph’s distances are left-skewed and bi-modal. In this specific insttance, the left mode reflects distances between nodes in the same blocks, whereas the right mode reflects distances between nodes not in the same blocks. Naturally, this bi-modal pattern does not appear in the case of the ER graph.
@@
Unfortunately, the K-S test requires that both the observations within each distribution and the distributions being compared themselves be independent and identically distributed (iid). While the $p$-values in our test do behave exactly as expected, they cannot be relied on for their statistical accuracy. For this reason, we focus our attention on the K-S distances themselves. After all, even with moderately sized graphs like the one in our experiments (e.g., $N=5,000\Rightarrow\sim 5,000^{2}$ distances), the sizes of our samples are extremely large. Argubaly, these extremely large samples make our distances very reliable metrics of the (dis)similarity between the distributions being compared. For this reason, under our simplified test, we consider two graph to be dissimilar, when the K-S distance is greater than zero, without regard for the significance of this figure.
@@ 4 Numerical results
We conduct four types of numerical experiments. In Section 4.1, we compare graphs that are known to be different. The goal of these comparisons is to illustrate our test. These comparisons are new instances of the tests we conducted in our previous work [18].
In Section 4.2, we compare individual graphs with versions of themselves with randomly selected nodes or edges having been removed. The goal of these experiments is, in part, to validate the correct verdict of similarity, in cases where we compare the graph to an unaltered version of itself. These trials also offer insight into the sensitivity of our test to graph structure modification through node or edge removal (e.g., in cases of temporal or change-point detection settings).
It is important to note here that these modifications to graphs (node or edge removal) may or may not constitute a significant or even observable modification to the initial graph. The experiments in this section consist of the same modification to graphs with varied structures. The goal of these experiments is not to set a threshold for two graphs to be declared “significantly dissimilar”. Rather, the goal is to examine if and how these modifications will be reflected in the K-S distances between the Jaccard distance distributions. Naturally, change-point or anomaly detection can be performed with our test, but it should be coupled with area expertise and a thorough examination of each graph’s structure.
Finally, our fourth set of experiments, shown in Section 4.2.3, is a comparison between graphs with an added number of nodes and randomly generated connections between them and their initial graph version. These experiments validate the applicability of our tests and its correct conclusions (similar or not), in the case of the very sparse graphs that were unsuited to our edge and node removal experiments in Section 4.2. Here again, these experiments aim to test the sensitivity or our metric. They consist of the same modification to graphs with varied structures. These modifications may or may not be significant, given each initial graph’s structure.
@@ 4.1 Two-graph tests
As stated earlier, the goal of these comparisons is to validate the correct rejection of the assumption of similarity of the two graphs being compared. We conduct 12 tests between synthetic graphs with known (dissimilar) structures. Table 1 shows cases where the structural differences between the graphs are arguably slight. The distance distributions for each of the graphs referenced in this table are presented in Figure 2 (Erdös-Rényi, ER) and Figure 3 (stochastic block model, SBM).
@@
All synthetic graphs in these experiments have the same number of vertices ($N=3,400$). The SBM graphs are all composed of 45 blocks of varying sizes ($n\in[50,99]$) but uniform in/out edge probabilities. These probabilities are shown in their naming convention. The ER graphs against which they are compared were generated using edge probabilities that are equal to the density of the SBM graphs in the comparisons. The CM graphs were generated using power law in/out degree sequences of exponent 3.5.
Table 1: Pairwise comparisons of ER graphs with given densities and SBM graphs with given within block/between block connection probabilities. (hard cases)
Graph 1 Graph 2 KS Statistic
ER 0.309 SBM 0.7/0.3 0.004
ER 0.309 (Directed) SBM 0.7/0.3 (Directed) 0.005
ER 0.213 SBM 0.8/0.2 0.013
ER 0.213 (Directed) SBM 0.8/0.2 (Directed) 0.018
ER 0.118 SBM 0.9/0.1 0.061
ER 0.118 (Directed) SBM 0.9/0.1 (Directed) 0.086
Table 2: Pairwise comparisons of CM graphs with a given number of nodes (3400), ER graphs with given densities, and SBM graphs with given within block/between block connection probabilities. (easier cases)
Graph 1 Graph 2 KS Statistic
CM3400 ER 0.213 0.999
CM3400 (Directed) ER 0.213 (Directed) 0.998
CM3400 SBM 0.9/0.1 0.999
CM3400 (Directed) SBM 0.9/0.1 (Directed) 0.998
SBM 0.9/0.1 ER 0.213 0.976
SBM 0.9/0.1 (Directed) ER 0.213 (Directed) 0.977
@@
Figure 2: Jaccard distance distributions for ER models with given densities.
@@
Figure 3: Jaccard distance distributions for stochastic block models with given within block/between block connection probabilities.
@@
Figure 4: Jaccard distance distributions for configuration models with 3400 nodes.
@@
4.2 Sensitivity analysis
The tests in this section have three different goals. First we compare each graph to itself. Here, we expect the hypothesis of similarity not be rejected. We also compare each graph to a version of itself with a portion of its nodes or edges removed. The goal of these trials is to examine the test’s sensitivity to various structural modifications. We compare every graph in this section to a version of itself with 0, 1, 5, 10, 25% of its (randomly selected) nodes or edges removed.
@@ 4.2.1 Synthetic graphs
We begin by removing a percentage of randomly selected nodes from each of our four ER graphs, and compare the distribution of Jaccard distances to the distribution of distances in the complete original graph. The K-S distances to the original graphs are presented in Table 3 through Table 10.
@@
Percent Removed 0 1 5 10 25
ER 0.5 0 0.003 0.008 0.013 0.038
ER 0.5 (Directed) 0 0.002 0.007 0.018 0.039
ER 0.333 0 0.003 0.008 0.016 0.038
ER 0.333 (Directed) 0 0.002 0.007 0.014 0.041
@@
Percent Removed 0 1 5 10 25
ER 0.5 0 0.189 0.759 0.979 1
ER 0.5 Directed 0 0.263 0.902 0.999 1
ER 0.333 0 0.104 0.481 0.799 0.998
ER 0.333 Directed 0 0.146 0.638 0.929 1
@@
Percent Removed 0 1 5 10 25
SBM 0.7/0.3 0 0.003 0.006 0.014 0.041
SBM 0.7/0.3 (Directed) 0 0.002 0.007 0.013 0.042
SBM 0.9/0.1 0 0.002 0.007 0.016 0.036
SBM 0.9/0.1 (Directed) 0 0.001 0.008 0.014 0.035
@@
Percent Removed 0 1 5 10 25
SBM 0.7/0.3 0 0.093 0.438 0.751 0.995
SBM 0.7/0.3 (Directed) 0 0.130 0.584 0.893 1
SBM 0.9/0.1 0 0.029 0.146 0.284 0.628
SBM 0.9/0.1 (Directed) 0 0.041 0.202 0.389 0.785
We complete our tests on synthetic graphs with Configuration Model graphs.
Table 7: Nodes removal experiment for two CM graphs
Percent Removed 0 1 5 10 25
CM 0 na na na na
CM (Directed) 0 0.000 0.000 0.001 na
Table 8: Edges removal experiment for two CM graphs
Percent Removed 0 1 5 10 25
CM 0 0.000 na na na
CM (Directed) 0 0.000 0.000 na na
Here, we note that both CM graphs quickly become disconnected, as we remove nodes or edges, which renders our test inapplicable. In order to better understand why our distance does not seem to capture node or edge removals, in the cases were the graph remains connected, we examine the densities of these graphs, to understand if and how these edge and node removals affect the overall graph structure.
Table 9: Densities under node removal experiments for two CM graphs
Percent Removed 0 1 5 10 25
CM 0.001 na na na na
CM (Directed) 0.001 0.001 0.001 0.001 na
Table 10: Densities under edge removal experiments for two CM graphs
Percent Removed 0 1 5 10 25
CM 0.001 0.001 na na na
CM (Directed) 0.001 0.001 0.001 na na
Our experiments removed a percentage of existing edges or nodes. Naturally, removing a percentage of a very small number of pre-existing edges cannot significantly alter the graph. The removal of a few nodes or edges only has a very minor effect on this reality. Therefore, the graph’s overall structure remains largely unaffected by our experiments and our test does not detect a meaningful change in the overall graph structure. Nevertheless, throughout the remainder of our experiments on synthetic graphs, our test produces results exactly as expected. The test detects dissimilarity as it increases and never classifies a graph as different from its unmodified self. We also note that, predictably, edge removal introduces a more marked variation in the graph than node removal.